OpenStudy (anonymous):

Please help!! Best answer gets a medal. A school graduation class wants to hire buses and vans for a trip to Jasper National Park. Each bus holds 40 students and 3 teachers and cost $1200 to rent. Each van holds 8 students and 1 teacher and costs $100 to rent. The school has at least 400 students wanting to go, but at most 36 teachers. What is the mininum transportation cost? Make vans equal "x"

OpenStudy (anonymous):

I guess it's Min Z= 100x+ 1200y st 8x+40y>= 400 3y+ x<= 36 and usual non negative constraints Am I right?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

I don't know if that's right. I don't understand the question :/

OpenStudy (anonymous):

I guess it's mininum z = 100x + 1200y 8x + 40y =400 3y + x<=36

OpenStudy (anonymous):

10 buses and 2 vans will transport all the students and teachers and there wold be extra room for all of the student's gear